600 people attended the premiere of a movie. Adult tickets $500 and children$200. Box office receipts totalled $2400,000, how many children attended the premiere?
Answer by mananth(16949)

let number of children be x
number of adults will be 600-x
..
child ticket = 200
adult ticket = 500
..
200x+500(600-x)=240,000
200x+300000-500x= 240,000
-300x=240000-300000
-300x=-60000
x= 200 children
I think the receipts should be 240, 000. I have worked on that basis